# Kida’s formula via Selmer complexes

### 13 Jan 2024 v1

@@ 1.1.3. Arithmetic applications
  • §4.1: The $p$-ramified Iwasawa modules for the cyclotomic $\mathbb{Z}_{p}$-extensions of totally real fields. The result is essentially equivalent to the original Kida’s formula by the Kummer duality, as will be discussed in §4.2.
  • §4.2: The minus components of split Iwasawa modules for the cyclotomic $\mathbb{Z}_{p}$-extensions of CM-fields. This is a variation of the original Kida’s formula, and indeed they are equivalent to each other.
  • §4.3: The $\Sigma$-ramified Iwasawa modules for ordinary CM-fields, where $\Sigma$ is a $p$-adic CM-type. By specializing to the case where the base field is an imaginary quadratic field, we recover a recent result of Murakami [17]. The general case seems to be new.
  • §4.4: The Selmer groups for ordinary elliptic curves. This recovers the result of Hachimori–Matsuno [6].
  • §4.5: The (multiply-)signed Selmer groups for supersingular elliptic curves. When the base field is $\mathbb{Q}$, this has an overlap with the work [21], but our formulation is more explicit and the general formulation should be new.
@@ 1.2. Review of previous work
The method of Iwasawa [8, pages 285–288] is adopted later by, e.g., Hachimori–Matsuno [6] and Murakami [17]. Firstly we reduce the problem to the special case where the extension is cyclic of degree $p$. For a cyclic group $G$ of order $p$, we have a classification theorem of $\mathbb{Z}_{p}[G]$-lattices. This reduces the task to the computation of Herbrand quotients of arithmetic modules.
@@
Finally, it is worth mentioning that Nichifor [20, Theorem 3.3] proved the projectivity of the Iwasawa modules by applying Kida’s formula. Conceptually, this is in the opposite direction to our perspective in this paper; we know the perfectness of arithmetic objects in advance, from which Kida’s formula follows.
@@ 2. The key algebraic theorem
Let $\Lambda$ be a ring that is isomorphic to $\mathbb{Z}_{p}[[T]]$ with $p$ a prime number. (The coefficient ring $\mathbb{Z}_{p}$ can be replaced by any $p$-adic integer ring, but to ease the notation we use $\mathbb{Z}_{p}$.) Let $G$ be a finite $p$-group.
@@ 2.1. Module version
This is a special case of [15, Proposition 2.4], which deals with more generally quasi-projective modules. For the convenience of the readers, we reproduce the proof, following [9, Proposition 8.5].
@@ 3.2. The Selmer complex
For the basic properties of these complexes, see the author’s paper [10, §3], or Fukaya–Kato [1, §1.6] and Nekovář [18] as more comprehensive references ([1] deals with non-commutative cases).
The Selmer complex $C$ in the current situation (see [18, §6.1]) is defined so that it satisfies a triangle
@@
    The complex $C$ is perfect over $\mathcal{R}_{K_{\infty}}$. More precisely, $C$ is quasi-isomorphic to a complex of the form $[C^{0}\to C^{1}\to C^{2}\to C^{3}]$ with finitely generated projective $\mathcal{R}_{K_{\infty}}$-modules $C^{i}$ located at degree $i$. This follows from the fact that the global and local Iwasawa cohomology complexes are perfect (see [1, Proposition 1.6.5(2)]).
@@
    over $\mathcal{R}_{k_{\infty}}$. This follows from the corresponding compatibility about the global and local Iwasawa cohomology complexes (see [1, Proposition 1.6.5(3)]).
@@
    The Euler characteristic of $C$ is zero. This follows from the global and local Euler–Poincare characteristic formulas and Assumption 3.1. Indeed, the Euler characteristic of $\operatorname{\mathsf{R}\Gamma}_{\operatorname{Iw}}(k_{S}/K_{\infty},T)$ (resp. $\bigoplus_{v\in S_{f}}\operatorname{\mathsf{R}\Gamma}_{\operatorname{Iw}}(K_{\infty,v},T_{v}^{-})$) is the right (resp. left) hand side of the equation in Assumption 3.1 (see [19, Theorems (7.3.1) and (8.7.4)]).
@@
    This follows from the complex version of the global duality (see [18, §5.4])
@@
By property (i), we have $H^{i}(C)=0$ for $i\neq 0,1,2,3$. It is known that both $H^{0}_{\operatorname{Iw}}(k_{S}/K_{\infty},T)$ and $H^{0}_{\operatorname{Iw}}(K_{\infty,v},T_{v}^{-})$ ($v\in S_{f}$) vanish by Assumption 3.2 (see Rubin [22, Appendix B, Lemma 3.2]). Then the definition of $C$ implies $H^{0}(C)=0$ and (3.5). On the other hand, property (iv) and the local duality induce (3.6) since the Selmer group $\operatorname{Sel}_{\mathcal{T}^{+}}(A/K_{\infty})$ coincides with the kernel of the natural homomorphism
@@
We note the following: If there exists a finite prime $v$ of $K$ such that $T_{v}^{-}=T$ (i.e., $T_{v}^{+}=0$), then we have $H^{1}(C)_{\operatorname{tors}}=0$, where $(-)_{\operatorname{tors}}$ denotes the maximal $\mathcal{R}_{K_{\infty}}$-torsion submodule. This easily follows by observing that we may take $C^{0}=0$ in property (i) in this case. In general, we have the following description of $H^{1}(C)_{\operatorname{tors}}$. The result should be known to experts (e.g., [18, Proposition 8.11.5(ii), Corollary 9.1.7]), but we include a direct proof for convenience.
@@ 4.1. Totally real fields
Let $K_{\infty}/k_{\infty}$ be a finite $p$-extension between the cyclotomic $\mathbb{Z}_{p}$-extensions of totally real fields. Let $X_{p}(K_{\infty})$ denote the $p$-ramified Iwasawa module for $K_{\infty}$ (that is, ramification at $p$-adic primes is allowed). The result is the following. It is already known by Iwasawa [8] (see also Neukirch–Schmidt–Wingberg [19, Corollary (11.4.11)]).
@@
    The neat formula in (2) relating $\lambda(X_{S}(K_{\infty}))$ and $\lambda(X_{S}(k_{\infty}))$ is not surprising. Indeed, taking the Kummer duality in mind, we see from [8, page 276, Theorem 3 and the succeeding text] that $\mu=0$ holds for $X_{p}(k_{\infty})$ if and only if $\operatorname{Gal}(k_{S}/k_{\infty})$ is a free pro-$p$-group. Then the formula is a consequence of the Schreier index formula on free groups (or one may apply [8, page 250, Lemma 7]).
@@ 4.2. Minus components for CM-fields
The original formula of Kida deals with $X(K_{\infty})^{-}$. However, it is not $X(K_{\infty})^{-}$ that Theorem 3.3 directly applies to, but $X^{\prime}(K_{\infty})^{-}$. The result is the following. It is already shown in [19, Corollary (11.4.13)], but the methods are different (see the end of this subsection).
@@
One is to use Theorem 4.1 and the Kummer duality (see [24, Propositions 13.32 and 15.37]) between the $p$-ramified Iwasawa modules for totally real fields and the minus components of the unramified Iwasawa modules for CM-fields. Strictly speaking, we need a slightly strengthened statement of Theorem 4.1, but we omit the details.
The other is to use Theorem 4.3 and the comparison between $X^{\prime}(K_{\infty})^{-}$ and $X(K_{\infty})^{-}$. For this, we only need to use [19, Lemma (11.4.9)], which gives us an exact sequence
@@ 4.4. Ordinary elliptic curves
Theorem 4.6 (Hachimori–Matsuno [6, Theorem 3.1]).
@@
where $\hat{E}_{v}$ denotes the associated formal group and $\tilde{E}_{v}$ the reduction modulo $v$. We also set $T_{v}^{+}=T$ for any $v\nmid p$. This data $\mathcal{T}^{+}$ satisfies Assumption 3.1. Then we have $\operatorname{Sel}(E/K_{\infty})=\operatorname{Sel}_{\mathcal{T}^{+}}(A/K_{\infty})$ (see Greenberg [2, Proposition 2.4]).
@@
  • (global) The group $E(K_{\infty})[p^{\infty}]$ is finite, thanks to Mazur [16, Proposition 6.12] (see [2, page 57]). This also implies $\delta_{\mathcal{T}^{+}}(T/K_{\infty})=0$.
@@
In [6], the prime $p=2$ is also dealt with when $k$ is totally imaginary. In principle our method can be applied to this case, but we omit it. On the other hand, we have removed an assumption when $p=3$ (denoted by (Hyp2) in [6]) that, for each non-$p$-adic prime, if we have additive reduction for $E/k_{\infty}$, then we still have additive reduction for $E/K_{\infty}$. This is why the “pot-good” or “pot-split” term does not appear in [6, Theorem 3.1].
@@ 4.5. Supersingular elliptic curves
Finally, let us sketch how to deduce Kida’s formula for supersingular elliptic curves. This demonstrates how Theorem 2.5 can also be applied to non-ordinary representations. Our result specialized to the case where the base field $F$ is $\mathbb{Q}$ should coincide with that of [21, Theorem 4.3] specialized to the case the modular form is associated to an elliptic curve, but our formulation is more explicit. When $F\neq\mathbb{Q}$, we need the author’s work [11] on the precise structures of $\pm$-local conditions.
@@
Set $F_{\infty}=F(\mu_{p^{\infty}})$. Let $K_{\infty}/F_{\infty}$ be a finite extension. We assume that $K_{\infty}/F$ is abelian in order to use results in [11]. Then as in [11, Definition 2.3], for each prime $\mathfrak{p}\in S_{p}(F)$ and for each sign $\pm$, we have a submodule $E^{\pm}(K_{\infty,\mathfrak{p}})$ of $E(K_{\infty,\mathfrak{p}})$.
Let us take a multi-sign $\epsilon=(\epsilon_{\mathfrak{p}})_{\mathfrak{p}}\in\prod_{\mathfrak{p}\in S_{p}(F)}\{+,-\}$. As in [11, Definition 2.5], we define the $\epsilon$-Selmer group $\operatorname{Sel}^{\epsilon}(E/K_{\infty})$ as the kernel of the natural map
@@
Then properties (i)–(iv) still hold. To show them, we only have to use [11, Proposition 4.4] on the module structure of $(E^{\pm}(K_{\infty,\mathfrak{p}})\otimes\mathbb{Q}_{p}/\mathbb{Z}_{p})^{\vee}$ (i.e., the finiteness of the projective dimension, the computation of the rank, and the compatibility with respect to the field extension).

### 01 Mar 2024 v2

@@ 1.1.3. Arithmetic applications
  • §4.1:
    The $p$-ramified Iwasawa modules for the cyclotomic $\mathbb{Z}_{p}$-extensions of totally real fields. The result is essentially equivalent to the original Kida’s formula by the Kummer duality, as will be discussed in §4.2.
  • §4.2:
    The minus components of split Iwasawa modules for the cyclotomic $\mathbb{Z}_{p}$-extensions of CM-fields. This is a variation of the original Kida’s formula, and indeed they are equivalent to each other.
  • §4.3:
    The $\Sigma$-ramified Iwasawa modules for ordinary CM-fields, where $\Sigma$ is a $p$-adic CM-type. By specializing to the case where the base field is an imaginary quadratic field, we recover a recent result of Murakami [21]. The general case seems to be new.
  • §4.4:
    The Selmer groups for ordinary elliptic curves. This recovers the result of Hachimori–Matsuno [9].
  • §4.5:
    The (multiply-)signed Selmer groups for supersingular elliptic curves. This basically recovers a result of Lim [19].
1.1.4. Analogues for $p$-adic Lie extensions
Finally in §5, we discuss Kida’s formula in the context of non-commutative Iwasawa theory, developed by Coates–Fukaya–Kato–Sujatha–Venjakob [2], among others. More precisely, we consider the case where $\operatorname{Gal}(K_{\infty}/k_{\infty})$ is a non-trivial, pro-$p$, $p$-adic Lie group without $p$-torsion (so in particular $K_{\infty}/k_{\infty}$ is infinite). We will show that our method established in §§24 can be also applied to this setting, which results in a formula on the $\mathbb{Z}_{p}[[\operatorname{Gal}(K_{\infty}/k_{\infty})]]$-rank of the Selmer group. Our result recovers a theorem of Hachimori–Sharifi [10] concerning the minus components of unramified Iwasawa modules for CM-fields.
@@ 1.2. Review of previous work
The method of Iwasawa [12, pages 285–288] is adopted later by, e.g., Hachimori–Matsuno [9] and Murakami [21]. Firstly we reduce the problem to the special case where the extension is cyclic of degree $p$. For a cyclic group $G$ of order $p$, we have a classification theorem of $\mathbb{Z}_{p}[G]$-lattices. This reduces the task to the computation of Herbrand quotients of arithmetic modules.
@@
As for $p$-adic Lie extensions, our method is still different from Hachimori–Sharifi [10] mentioned in §1.1.4. They determined the rank of the minus component of the unramified Iwasawa module by applying the original Kida’s formula to the finite intermediate extensions and then investigating the asymptotic growth pattern of the $\lambda$-invariants. Our method is more direct.
Finally, it is worth mentioning that Nichifor [24, Theorem 3.3] proved the projectivity of the Iwasawa modules by applying Kida’s formula. Conceptually, this is in the opposite direction to our perspective in this paper; we know the perfectness of arithmetic objects in advance, from which Kida’s formula follows.
Acknowledgments
@@ 2. The key algebraic theorem
Let $\Lambda$ be a ring that is isomorphic to $\mathbb{Z}_{p}[[T]]$ with $p$ a prime number. (The coefficient ring $\mathbb{Z}_{p}$ can be replaced by any $p$-adic integer ring, but to ease the notation we use $\mathbb{Z}_{p}$.) Let $G$ be a finite $p$-group. Throughout this paper we always mean left modules by simply saying modules.
@@ 2.1. Module version
This is a special case of [19, Proposition 2.4], which deals with more generally quasi-projective modules. For the convenience of the readers, we reproduce the proof, following [13, Proposition 8.5].
@@ 3.2. The Selmer complex
For the basic properties of these complexes, see the author’s paper [14, §3], or Fukaya–Kato [4, §1.6] and Nekovář [22] as more comprehensive references ([4] deals with non-commutative cases).
The Selmer complex $C$ in the current situation (see [22, §6.1]) is defined so that it satisfies a triangle
@@
    The complex $C$ is perfect over $\mathcal{R}_{K_{\infty}}$. More precisely, $C$ is quasi-isomorphic to a complex of the form $[C^{0}\to C^{1}\to C^{2}\to C^{3}]$ with finitely generated projective $\mathcal{R}_{K_{\infty}}$-modules $C^{i}$ located at degree $i$. This follows from the fact that the global and local Iwasawa cohomology complexes are perfect (see [4, Proposition 1.6.5(2)]).
@@
    over $\mathcal{R}_{k_{\infty}}$. This follows from the corresponding compatibility about the global and local Iwasawa cohomology complexes (see [4, Proposition 1.6.5(3)]).
@@
    The Euler characteristic of $C$ is zero. This follows from the global and local Euler–Poincare characteristic formulas and Assumption 3.1. Indeed, the Euler characteristic of $\operatorname{\mathsf{R}\Gamma}_{\operatorname{Iw}}(k_{S}/K_{\infty},T)$ (resp. $\bigoplus_{v\in S_{f}}\operatorname{\mathsf{R}\Gamma}_{\operatorname{Iw}}(K_{\infty,v},T_{v}^{-})$) is the right (resp. left) hand side of the equation in Assumption 3.1 (see [23, Theorems (7.3.1) and (8.7.4)]).
@@
    This follows from the complex version of the global duality (see [22, §5.4])
@@
By property (i), we have $H^{i}(C)=0$ for $i\neq 0,1,2,3$. It is known that both $H^{0}_{\operatorname{Iw}}(k_{S}/K_{\infty},T)$ and $H^{0}_{\operatorname{Iw}}(K_{\infty,v},T_{v}^{-})$ ($v\in S_{f}$) vanish by Assumption 3.2 (see Rubin [26, Appendix B, Lemma 3.2]). Then the definition of $C$ implies $H^{0}(C)=0$ and (3.5). On the other hand, property (iv) and the local duality induce (3.6) since the Selmer group $\operatorname{Sel}_{\mathcal{T}^{+}}(A/K_{\infty})$ coincides with the kernel of the natural homomorphism
@@
We note the following: If there exists a finite prime $v$ of $K$ such that $T_{v}^{-}=T$ (i.e., $T_{v}^{+}=0$), then we have $H^{1}(C)_{\operatorname{tors}}=0$, where $(-)_{\operatorname{tors}}$ denotes the maximal $\mathcal{R}_{K_{\infty}}$-torsion submodule. This easily follows by observing that we may take $C^{0}=0$ in property (i) in this case. In general, we have the following description of $H^{1}(C)_{\operatorname{tors}}$. The result should be known to experts (e.g., [22, Proposition 8.11.5(ii), Corollary 9.1.7]), but we include a direct proof for convenience.
@@ 4.1. Totally real fields
Let $K_{\infty}/k_{\infty}$ be a finite $p$-extension between the cyclotomic $\mathbb{Z}_{p}$-extensions of totally real fields. Let $X_{p}(K_{\infty})$ denote the $p$-ramified Iwasawa module for $K_{\infty}$ (that is, ramification at $p$-adic primes is allowed). The result is the following. It is already known by Iwasawa [12] (see also Neukirch–Schmidt–Wingberg [23, Corollary (11.4.11)]).
@@
    The neat formula in (2) relating $\lambda(X_{S}(K_{\infty}))$ and $\lambda(X_{S}(k_{\infty}))$ is not surprising. Indeed, taking the Kummer duality in mind, we see from [12, page 276, Theorem 3 and the succeeding text] that $\mu=0$ holds for $X_{p}(k_{\infty})$ if and only if $\operatorname{Gal}(k_{S}/k_{\infty})$ is a free pro-$p$-group. Then the formula is a consequence of the Schreier index formula on free groups (or one may apply [12, page 250, Lemma 7]).
@@ 4.2. Minus components for CM-fields
The original formula of Kida deals with $X(K_{\infty})^{-}$. However, it is not $X(K_{\infty})^{-}$ that Theorem 3.3 directly applies to, but $X^{\prime}(K_{\infty})^{-}$. The result is the following. It is already shown in [23, Corollary (11.4.13)], but the methods are different (see the end of this subsection).
@@
One is to use Theorem 4.1 and the Kummer duality (see [28, Propositions 13.32 and 15.37]) between the $p$-ramified Iwasawa modules for totally real fields and the minus components of the unramified Iwasawa modules for CM-fields. Strictly speaking, we need a slightly strengthened statement of Theorem 4.1, but we omit the details.
The other is to use Theorem 4.3 and the comparison between $X^{\prime}(K_{\infty})^{-}$ and $X(K_{\infty})^{-}$. For this, we only need to use [23, Lemma (11.4.9)], which gives us an exact sequence
@@ 4.4. Ordinary elliptic curves
Theorem 4.6 (Hachimori–Matsuno [9, Theorem 3.1]).
@@
where $\hat{E}_{v}$ denotes the associated formal group and $\tilde{E}_{v}$ the reduction modulo $v$. We also set $T_{v}^{+}=T$ for any $v\nmid p$. This data $\mathcal{T}^{+}$ satisfies Assumption 3.1. Then we have $\operatorname{Sel}(E/K_{\infty})=\operatorname{Sel}_{\mathcal{T}^{+}}(A/K_{\infty})$ (see Greenberg [5, Proposition 2.4]).
@@
  • (global) The group $E(K_{\infty})[p^{\infty}]$ is finite, thanks to Mazur [20, Proposition 6.12] (see [5, page 57]). This also implies $\delta_{\mathcal{T}^{+}}(T/K_{\infty})=0$.
@@
In [9], the prime $p=2$ is also dealt with when $k$ is totally imaginary. In principle our method can be applied to this case, but we omit it. On the other hand, we have removed an assumption when $p=3$ (denoted by (Hyp2) in [9]) that, for each non-$p$-adic prime, if we have additive reduction for $E/k_{\infty}$, then we still have additive reduction for $E/K_{\infty}$. This is why the “pot-good” or “pot-split” term does not appear in [9, Theorem 3.1].
@@ 4.5. Supersingular elliptic curves
As a final application, let us sketch how to deduce Kida’s formula for supersingular elliptic curves. This demonstrates how Theorem 2.5 can also be applied to non-ordinary representations. Our result is basically the same as [19, Proposition 5.2] (with minor differences such as Remark 4.8). Also, when the base field $F$ is $\mathbb{Q}$, the result should coincide with that of [25, Theorem 4.3] specialized to the case the modular form is associated to an elliptic curve, but our formulation is more explicit. When $F\neq\mathbb{Q}$, we need the author’s work [15] on the precise structures of $\pm$-local conditions.
@@
Set $F_{\infty}=F(\mu_{p^{\infty}})$. Let $K_{\infty}/F_{\infty}$ be a finite extension. We assume that $K_{\infty}/F$ is abelian in order to use results in [15]. Then as in [15, Definition 2.3], for each prime $\mathfrak{p}\in S_{p}(F)$ and for each sign $\pm$, we have a submodule $E^{\pm}(K_{\infty,\mathfrak{p}})$ of $E(K_{\infty,\mathfrak{p}})$.
Let us take a multi-sign $\epsilon=(\epsilon_{\mathfrak{p}})_{\mathfrak{p}}\in\prod_{\mathfrak{p}\in S_{p}(F)}\{+,-\}$. As in [15, Definition 2.5], we define the $\epsilon$-Selmer group $\operatorname{Sel}^{\epsilon}(E/K_{\infty})$ as the kernel of the natural map
@@
Then properties (i)–(iv) still hold. To show them, we only have to use [15, Proposition 4.4] on the module structure of $(E^{\pm}(K_{\infty,\mathfrak{p}})\otimes\mathbb{Q}_{p}/\mathbb{Z}_{p})^{\vee}$ (i.e., the finiteness of the projective dimension, the computation of the rank, and the compatibility with respect to the field extension).